# GeoFaaS: An Edge-to-Cloud FaaS Platform Short Paper

### v1

@@ I Introduction
This paper addresses the limitation of inherent geographical context awareness in FaaS platforms. We propose GeoFaaS (Geo-aware Function-as-a-Service), a geographically distributed platform operating across the heterogeneous edge-to-cloud continuum. By leveraging client location information, GeoFaaS transparently routes function requests to the nearest available FaaS resource, minimizing network latency for improved performance. Additionally, GeoFaaS supports transparent offloading by means of fallback processing in the cloud to handle edge failures resulting from, e.g., overload situations. Building on the single-node FaaS system tinyFaaS [3], GeoFaaS works as a stand-alone edge-to-cloud FaaS platform. Nevertheless, GeoFaaS can use its modular architecture to integrate existing FaaS services, e.g., in the cloud, and also act as a routing layer only.
@@ III GeoFaaS Architecture
Figure 1: The GeoFaaS system with two GeoFaaS nodes deployed on edge and one node deployed on the cloud. The Client at time $t_{0}$ is outside of edge service areas, so it communicates with cloud. When client moves to a new service area at $t_{1}$, it will be passed to a new node for further communication by its last responsible node.
Figure 2: The GeoFaaS node architecture comprises three key elements: GeoBroker, which enables location-aware pub/sub messaging, a local FaaS server or a FaaS service running on external servers within the same data center, and the GeoFaaS Bridge, a middleware between the other elements.
@@ III-A Distributed Geo-Context Aware Message Broker
In order for a client to initiate communication with the system, it has the freedom to target any GeoBroker instance (no matter where, possibly cached from a previous request or fetched from DNS) to publish its first request. Then DisGB acts as a smart intermediary, directing the request to the geographically closest GeoFaaS Bridge (subscriber) to the client. From this point, for any follow-up requests, the client will communicate directly with its responsible GeoBroker instance. Note that, disjoint edge broker areas are complemented by the cloud’s broker area, guaranteeing total coverage. This means that even when a client transitions beyond the boundaries of its current GeoBroker’s area, as depicted in Fig. 1, the current GeoBroker will pass the client to the new GeoBroker. This ensures that each client is connected to the geographically most relevant GeoBroker instance for further communication.
@@ III-C GeoFaaS Bridge
As mentioned before, each GeoFaaS node has a predefined service area which corresponds to the DisGB broker areas. The GeoFaaS Bridge uses this service area as a filter criterion for subscribing to function calls, ensuring only relevant requests are received. While edge GeoFaaS nodes mirror their corresponding GeoBroker’s broker area as their service area, the service area of the cloud covers the entire world.
Each GeoFaaS Bridge and its associated GeoBroker instance operate autonomously, making independent decisions within their respective service areas.
@@ IV-A Proof-of-Concept Implementation
We implement the GeoFaaS Bridge and the client library in Kotlin and use tinyFaaS’ HTTP endpoint. For each deployed function, we create five topics in DisGB. In case of an $f1$ function, the specific topics used are detailed in Table I. Upon calling a function via /call, the GeoFaaS client actively listens for both the /ack and /result topics associated with the function, ensuring it receives both an acknowledgment of its request and the function’s result. The cloud is responsible for the requests outside all (edge) service areas, therefore its subscription geofence is the world. Unlike the cloud, the edge subscribes only to /call topics per function. As indicated in Table I, for each function, the cloud subscribes to two additional topics: one for direct client calls in case of unresponsive edges (retry by cloud), and another for edge nodes signaling their inability to serve a request. In the following experiments, we deactivate “retry by cloud”, as we expect results to be delivered without needing the /call/retry topic for direct cloud calls.
@@ IV-B1 Distance/Latency Change Scenario
Figure 4: Each client’s movement is followed by a function call, with response latencies plotted in blue (ms). The red line shows client-to-server distance (km), and vertical lines indicate switching to a new GeoFaaS node based on location updates.
@@ IV-B2 High Load Scenario
This section demonstrates the ability of GeoFaaS to maintain continuous service transparently, ensuring clients do not need to intervene when requests are offloaded by GeoFaaS. As the focus of this scenario is testing the offloading functionality, this experiment comprises just one edge and one cloud server. To this end, we employ a computationally intensive function to push the edge server to its limits, thereby increasing the likelihood of requests being offloaded to the cloud for processing. The function difficulty is increased to calculate prime numbers under 1 million (instead of 10k) on both servers to further overload the edge server. We also set the client’s “result timeouts” to be high to prevent it from retrying while its request is being processed.
@@ V Conclusion and Future Work
This work addressed the growing demand for geographically distributed computing by proposing GeoFaaS, a novel edge-to-cloud FaaS platform. GeoFaaS leverages real-time client location information to transparently route function calls to the nearest available FaaS resource, minimizing network hops and latency. Additionally, it showed seamless execution with a cloud node, even with limited edge resources. We developed and evaluated an open-source proof-of-concept prototype, demonstrating the feasibility and effectiveness of GeoFaaS.
In future work, we aim to focus on optimizing function placement to ensure functions are deployed in locations that minimize latency and maximize resource utilization. Additionally, we will explore and experiment optimized policies for FaaS server selection or cloud offloading based on factors such as utilization and resource availability.

### v2

@@ I Introduction
This paper addresses the limitation of inherent geographical context awareness in FaaS platforms. We propose GeoFaaS (Geo-aware Function-as-a-Service), a geographically distributed FaaS platform operating across the heterogeneous edge-to-cloud continuum. By leveraging client location information, GeoFaaS transparently routes function requests to the nearest available FaaS resource, minimizing network latency for improved performance. Additionally, GeoFaaS supports transparent offloading by means of fallback processing in the cloud to handle edge failures resulting from, e.g., overload situations. Building on the single-node FaaS system tinyFaaS [3], GeoFaaS works as a stand-alone edge-to-cloud FaaS platform. Nevertheless, GeoFaaS can use its modular architecture to integrate existing FaaS services, e.g., in the cloud, and also act as a routing layer only.
@@ III GeoFaaS Architecture
Figure 1: The GeoFaaS system with two GeoFaaS nodes deployed on edge and one node deployed on the cloud. The Client at time $t_{0}$ is outside of edge service areas, so it communicates with cloud. When client moves to a new service area at $t_{1}$, the cloud node will hand off the client to a new node for further communication by its last responsible node.
Figure 2: The GeoFaaS node architecture comprises three key components: GeoBroker, which enables location-aware pub/sub messaging, a local FaaS server and/or FaaS services running on external servers within the same data center, and the GeoFaaS Bridge, a middleware between the other components.
@@ III-A Distributed Geo-Context Aware Message Broker
In order for a client to initiate communication with the system, it has the freedom to target any GeoBroker instance (no matter where, possibly cached from a previous request or fetched from DNS) to publish its first request. Then DisGB acts as a smart intermediary, directing the request to the geographically closest GeoFaaS Bridge (subscriber) to the client. From this point, for any follow-up requests, the client will communicate directly with its responsible GeoBroker instance. Note that, disjoint edge broker areas are complemented by the cloud’s broker area, guaranteeing total coverage. This means that even when a client transitions beyond the boundaries of its current GeoBroker’s area, as depicted in Fig. 1, the current GeoBroker will hand off the client to the new responsible GeoBroker. This ensures that each client is connected to the geographically most relevant GeoBroker instance for further communication.
@@ III-C GeoFaaS Bridge
As mentioned before, each GeoFaaS node has a predefined service area which corresponds to the DisGB broker areas. The GeoFaaS Bridge uses this service area as a filter criterion for subscribing to function calls, ensuring only relevant requests are received. While edge GeoFaaS nodes mirror their corresponding GeoBroker’s broker area as their service area, the service area of the cloud covers the entire world. Each GeoFaaS Bridge and its associated GeoBroker instance operate autonomously, making independent decisions within their respective service areas.
@@ IV-A Proof-of-Concept Implementation
We implement the GeoFaaS Bridge and the client library in Kotlin and use the HTTP endpoint of tinyFaaS. For each deployed function, we create five topics in DisGB. In case of an $f1$ function, the specific topics used are detailed in Table I. Upon calling a function via /call, the GeoFaaS client actively listens for both the /ack and /result topics associated with the function, ensuring it receives both an acknowledgment of its request and the function’s result. The cloud is responsible for the requests outside all (edge) service areas, therefore its subscription geofence is the world. Unlike the cloud, the edge subscribes only to /call topics per function. As indicated in Table I, for each function, the cloud subscribes to two additional topics: one for direct client calls in case of unresponsive edges (retry by cloud), and another for edge nodes signaling their inability to serve a request. In the following experiments, we deactivate “retry by cloud”, as we expect results to be delivered without needing the /call/retry topic for direct cloud calls.
@@ IV-B1 Distance/Latency Change Scenario
Figure 4: Each client’s movement is followed by a function call, with response latencies plotted in blue (ms). The red line shows client-to-server distance (km), and vertical lines indicate switching to a new GeoFaaS node based on location updates. The client’s movement is illustrated in Fig. 3.
@@ IV-B2 High Load Scenario
This section demonstrates the ability of GeoFaaS to maintain continuous service transparently, ensuring clients do not need to intervene when requests are offloaded by GeoFaaS. As the focus of this scenario is testing the offloading functionality, this experiment comprises just one edge and one cloud node. To this end, we employ a computationally intensive function to push the edge server to its limits, thereby increasing the likelihood of requests being offloaded to the cloud for processing. The function difficulty is increased by a factor of 100 to calculate prime numbers under 1 million on both servers to further overloading the edge server. We also set the clients’ “result timeouts” to be high to prevent them from retrying while their request is being processed.
@@ V Conclusion and Future Work
This work addressed the growing demand for geographically distributed computing by proposing GeoFaaS, a novel edge-to-cloud FaaS platform that prioritizes abstraction. GeoFaaS leverages real-time client location information to transparently route function calls to the nearest available FaaS resource, minimizing network hops and latency. By abstracting away detailed handover mechanisms, GeoFaaS provides a transparent execution experience, even with limited edge resources. We developed and evaluated an open-source proof-of-concept prototype, demonstrating the feasibility and effectiveness of GeoFaaS.
In future work, we aim to optimize function placement to ensure functions are deployed in locations that minimize latency and maximize resource utilization. Additionally, we will explore and experiment with optimized policies for FaaS server selection and cloud offloading based on factors such as utilization and resource availability.